# Stability and bifurcation analysis in a mechanochemical model of pattern formation

### v1

@@ 1. Introduction
By deriving a reduced one-dimensional model that preserves the essential mechanochemical feedback while simplifying the geometric setting (Section 3), we obtain, for the first time, a framework amenable to rigorous analytical investigation of the dynamics. We establish the existence of spatially heterogeneous steady states for sufficiently small diffusion coefficients and prove uniqueness of the homogeneous equilibrium for large diffusion, deriving explicit bounds on the corresponding critical values by variational methods (Section 4). This analysis explains the mechanism of symmetry breaking, demonstrating how patterned states bifurcate from the homogeneous configuration. We then investigate the pattern selection problem (Section 5) and show that only unimodal stationary patterns are linearly stable, while all multimodal solutions are unstable, revealing a fundamental distinction from classical Turing systems and curvature-based mechanochemical models [38, 31]. Finally, we identify both subcritical and supercritical pitchfork bifurcations and reveal additional fold bifurcations responsible for the emergence of bistable regimes (Section 6). Throughout, the theoretical results are complemented by numerical simulations, elucidating the interplay between nonlocality, mechanical feedback, and nonlinear kinetics, and providing a mathematically grounded framework for experimental exploration.
@@ 2. Main results
We start by formulating main theoretical results of this paper. We focus the mechano-chemical pattern formation model describing the concentration of the diffusive signaling molecules $u:=u(x,t)$, given by
@@
with the coefficients $D>0$ and $\kappa>0$, supplemented with the periodic boundary conditions $u(0,t)=u(1,t)$ and $\partial_{x}u(0,t)=\partial_{x}u(1,t)$ for all $t\geqslant 0$. This model arises from a one-dimensional formulation of a hyperelastic material governed by the Saint-Venant–Kirchhoff constitutive law. A detailed derivation is provided in Section 3. Numerical simulations of the model (Figure fig:numerics) demonstrate its capacity to produce stable unimodal patterns.
@@
    $\Gamma_{n}(s)=\big(\varphi_{n}(s),\kappa_{n}(s)\big),$
    admitting the asymptotic expansion
    (2.8) $\varphi_{n}(s)$ $=\kappa_{n}+s\sqrt{2}\cos(2\pi nx)+\mathcal{O}(s^{2}),\ \kappa_{n}(s)$ $=\kappa_{n}+\alpha_{n}^{\prime\prime}(0)s^{2}+\mathcal{O}(s^{3}),$
    as $s\to 0$.
@@ 3. Model derivation
Figure 3.1. Schematic representation of the elastic deformation framework for the Hydra spheroid model. Left: Reference configuration $S_{0}$ parametrized by $g:[0,L]\to S_{0}$. Right: Deformed configuration $S$ showing the composition of mappings $f=h\circ g$ and the arc-length reparametrization $\tilde{f}$ with $f=\tilde{f}\circ s$.
3.1. Hyper-elastic tissue model
@@
Building on the assumptions of the mechanochemical patterning model for Hydra tissue spheroids introduced in [41], we derive a reduced one-dimensional formulation that retains the essential coupling between mechanics and chemistry while simplifying the geometric structure. In this setting, the Hydra spheroid is represented as an elastic closed curve, providing a minimal framework to analyze the core mechano-chemical feedback mechanisms, see Figure 3.1.
(3.1) $g(\xi)=\frac{L}{2\pi}\begin{pmatrix}\cos\frac{2\pi}{L}\xi\\ \sin\frac{2\pi}{L}\xi\end{pmatrix},\quad\xi\in[0,L].$
This reference curve is subjected to a spatial deformation $h:\mathcal{S}_{0}\to\mathcal{S}$, resulting in a new closed curve $\mathcal{S}$ with circumference $\ell>0$. We prescribe a fixed enclosed area $A_{0}>\operatorname{Area}(\mathcal{S}_{0})=\pi L^{2}$ and restrict attention to deformations that satisfy $\operatorname{Area}(\mathcal{S})=A_{0}$. The deformed curve is parametrized by the composition
(3.2) $f:[0,L]\to\mathcal{S},\quad f=h\circ g.$
(3.3) $f(\xi)=\tilde{f}(s(\xi)),$
where $\tilde{f}:[0,\ell]\to\mathcal{S}$ satisfies $|\tilde{f}^{\prime}|=1$, and $s:[0,L]\to[0,\ell]$ encodes the tangential deformation. The deformation gradient in surface coordinates is
(3.4) $F=\nabla_{X}h(X)=f^{\prime}(\xi)\otimes\nabla\big(g^{-1}(X)\big),\quad X=g(\xi),$
and since $\nabla(g^{-1}(X))=g^{\prime}(\xi)$, we have
(3.5) $F=f^{\prime}(\xi)g^{\prime}(\xi)^{T}.$
The Cauchy (Right Cauchy-Green) strain tensor reads
(3.6) $C=F^{T}F=g^{\prime}f^{\prime T}f^{\prime}g^{\prime T},$
and the Green-Lagrangian strain is
(3.7) $\varepsilon=\frac{1}{2}(C-I)=\frac{1}{2}\big(|f^{\prime}|^{2}g^{\prime}g^{\prime T}-I\big).$
(3.8) $W(\varepsilon)=\frac{\lambda}{2}{\rm tr}^{2}(\varepsilon)+\frac{\mu}{2}{\rm tr}(\varepsilon^{2}),$
where the Lamé parameters satisfy $\lambda=0$ and $\mu=E(u)$, with $u:=u(\xi)$ the morphogen concentration and $E(u)$ the morphogen-dependent elastic modulus encoding the mechanochemical coupling central to our model. The specific choice of parameters follows from the analysis in [41]. The total elastic energy of a deformation is given by
(3.9) $\mathcal{E}_{\text{el}}=\int_{S_{0}}W(\varepsilon)dS_{0}=\frac{1}{8}\int_{0}^{L}{E(u)}\Big(\left(|f^{\prime}|^{2}-1\right)^{2}+1\Big)d\xi,$
where we used equation (3.7) and the identity $\operatorname{tr}(g^{\prime}\cdot g^{\prime T})=|g^{\prime}(\xi)|^{2}=1$ and $\operatorname{tr}I=2$.
(3.10) $A(f)=\int_{S}\operatorname{n}dS=\int_{0}^{L}f\cdot\operatorname{n}(f)|f^{\prime}|\ {d}\xi$
of the shape is preserved, i.e. $A(f)=A_{0}$. To find minimizers of $\mathcal{E}_{\text{el}}$ we introduce Lagrange multipliers. The Gateaux derivative with respect to variations $\delta f(\xi)$ with periodic boundary conditions yields
(3.11) $D\mathcal{E}_{\text{el}}(f)(\delta f)+pD(A-A_{0})(f)(\delta f)$ $=\frac{1}{2}\int_{0}^{L}E(u)\Big(|f^{\prime}|^{2}-1\Big)f^{\prime}\cdot\delta f^{\prime}\ d\xi$
$\quad+p\int_{0}^{L}\operatorname{n}(f)|f^{\prime}|\cdot\delta f\ d\xi=0,$
where $p$ is Lagrange multiplier and $n(x)$ denotes the outer unit-normal vector at $x$. Integrating by parts, we obtain that equation (3.11) is satisfied if and only if the minimizer $f$ satisfies
(3.12) $\frac{1}{2}\frac{d}{d\xi}\Big(E(u)\big(|f^{\prime}|^{2}-1\big)f^{\prime}\Big)+p\cdot\operatorname{n}(f)|f^{\prime}|=0.$
Substituting $f^{\prime}(\xi)=\tilde{f}^{\prime}(s(\xi))s^{\prime}(\xi)$ into equation (3.12), setting $\tilde{\varepsilon}=\frac{1}{2}\big(s^{\prime 2}-1\big)$ and differentiating the product, we obtain
(3.13) $0=\frac{d}{d\xi}\Big(E(u)\>\tilde{\varepsilon}\>s^{\prime}\Big)\tilde{f}^{\prime}(s)+E(u)\>\tilde{\varepsilon}\>s^{\prime}\>\tilde{f}^{\prime\prime}(s)s^{\prime}+p\cdot\operatorname{n}(f)s^{\prime}.$
Taking the scalar product of (3.13) with $\tilde{f}^{\prime}(s(\xi))$ and using $2\tilde{f}^{\prime\prime}\cdot\tilde{f}^{\prime}=\left(|\tilde{f}^{\prime}|^{2}\right)^{\prime}=0$ together with $\operatorname{n}(f)\cdot\tilde{f}^{\prime}=0$ yields
(3.14) $0=\frac{d}{d\xi}\Big(E(u)\>\tilde{\varepsilon}\>s^{\prime}\Big),$
and hence $E\big(u(\xi)\big)\tilde{\varepsilon}(\xi)s^{\prime}(\xi)$ is constant. Moreover, taking the scalar product with the rescaled outer unit-normal vector $s^{\prime}(\xi)^{-1}\operatorname{n}(f)$ on $\mathcal{S}$, we obtain
(3.15) $0=E(u)\tilde{\varepsilon}s^{\prime}\tilde{f}^{\prime\prime}(s)\cdot\operatorname{n}(f)+p$
yielding that the curvature $\tilde{f}^{\prime\prime}(s)\cdot\operatorname{n}(f)$ of $\mathcal{S}$ is constant along $\mathcal{S}$. Since the area enclosed by $\mathcal{S}$ is fixed, the minimizer $\mathcal{S}$ must be circular with circumference $\ell>L$. We can therefore fix the arc-length parametrization $\tilde{f}$ of $\mathcal{S}$ even if $E(u)$ changes with time and describe the deformation solely through the mapping $s:[0,L]\to[0,\ell]$, where $f=\tilde{f}\circ s$. Since $s^{\prime}(\xi)$ represents the local stretch ratio (ratio of deformed to reference arc length), the quantity
(3.16) $\tilde{\varepsilon}=\frac{1}{2}(s^{\prime 2}-1)$
is the finite Green-Lagrangian strain in one dimension.
Conservation of total length imposes
(3.17) $\int_{0}^{L}(s^{\prime}(\xi)-1)\,d\xi=\ell-L.$
The morphogen concentration $u(t,\xi)$ evolves according to
(3.18) $\partial_{t}u=\kappa\tilde{\varepsilon}-\alpha u+D\partial_{\xi\xi}u,$
with production proportional to local strain, linear degradation rate $\alpha>0$, and diffusion coefficient $D>0$. Taking together, the full mechano-chemical system reads
(3.19) $\begin{aligned} \partial_{t}u&=\kappa\tilde{\varepsilon}-\alpha u+D\partial_{\xi\xi}u,\\ 0&=\partial_{\xi}\big(E(u)(s^{\prime 2}-1)s^{\prime}\big),\end{aligned}$ $\xi\in[0,L],\ t\geq 0,$
$\ell-L=\int_{0}^{L}(s^{\prime}-1)\,d\xi,$ $t\geq 0,$
with periodic boundary conditions and appropriate initial data. The 1D finite strain is defined by (3.16), completing the mechanochemical coupling.
3.2. Inifinitesimal strain
To facilitate analytical treatment of the mechanochemical model, we invoke the infinitesimal strain approximation. This is justified biologically, as Hydra epithelial cells experience only small-amplitude displacements during mechanochemical signaling, so that the local deformation gradients remain close to the identity [41]. Under the assumption that $|s^{\prime}-1|\ll 1,$ the finite strain $\tilde{\varepsilon}=\tfrac{1}{2}(s^{\prime 2}-1)$ can be replaced by its linearized counterpart
(3.20) $\hat{\varepsilon}(t,\xi)=s^{\prime}(t,\xi)-1\approx\tilde{\varepsilon}(t,\xi),$
[…]
@@ 3.2. Inifinitesimal strain
(3.22) $\begin{aligned} \partial_{t}u&=\kappa\,\varepsilon-\alpha u+D\,\partial_{\xi\xi}u,\\ 0&=\partial_{\xi}\big(E(u)\,\varepsilon\big),\end{aligned}$ $\xi\in[0,L],\quad t\geq 0,$
$l-L=\int_{0}^{L}\varepsilon\,d\xi,$ $t\geq 0.$
@@
The third equation in (LABEL:sys_eps) expresses conservation of total strain, effectively introducing mechanical long-range inhibition, which acts alongside the positive feedback between local strain and morphogen concentration. This combination implements the classical local activation - long-range inhibition (LALI) mechanism.
@@
From the momentum balance equation in (LABEL:sys_eps), the stress depends only on time $t$, that is,
@@
(3.27) $\varepsilon(t,\xi)=\frac{\sigma(t)}{E(u(t,\xi))}=\frac{l-L}{E(u(t,\xi))\int_{0}^{L}\frac{1}{E(u(t,x))}\,dx}.$
Using the closed-form expression for the strain, the morphogen equation in (LABEL:sys_eps) can be written as
(3.28) $\partial_{t}u(t,\xi)=\frac{\kappa(l-L)}{E(u(t,\xi))\int_{0}^{L}\frac{1}{E(u(t,x))}\,dx}-\alpha u+D\,\partial_{\xi\xi}u\,.$
@@ 4.1. Variational formulation
There exist $\{f_{k}\}_{k=0}^{\infty}$ an orthonormal basis of $L^{2}(\mathbb{T})$ and a orthogonal basis of $W^{1,2}(\mathbb{T})$ and a set of eigenvalues $\{\mu_{k}\}_{k=0}^{\infty}\subseteq[0,\;\infty)$, such that $(f_{k},\mu_{k})$ is the eigenpair of the operator $\Delta$
@@ 5.3. Relation between eigenvalues
It is required to take care of the terms with $\beta_{n}=0$ as they vanish in the series (5.8). These sets are not necessarily disjoint, i.e. there can exists a double eigenvalue $\lambda_{n}=\lambda_{n+1}$ and $\lambda_{n}\in\Lambda_{0}$ while $\lambda_{n+1}\in\Lambda_{\beta}$. Without loss of generality we assume that the elements of $\Lambda_{0},\Lambda_{\beta}$ are in monotonically decreasing order
@@ 5.4. Stability of steady states
which is exactly the local problem (5.6) with $\lambda=0$. Since $U\in W^{2,2}(\mathbb{T})$, all $U$, $U_{x}$ and $U_{xx}$ satisfy periodic boundary conditions. Hence, the local problem has a zero eigenvalue with corresponding eigenfunction with at least $3$ zeros on the interval $[0,1)$. Proposition 5.3 implies that $\lambda_{3}\geqslant 0$ as the eigenfunctions $\psi_{0}$, $\psi_{1}$, $\psi_{2}$ have at most two zeros on the interval $[0,1)$. Therefore there are at least three strictly positive local eigenvalues $\lambda_{0}>\lambda_{1}\geqslant\lambda_{2}>\lambda_{3}\geq 0$. This is already sufficient to obtain a positive nonlocal eigenvalue $\nu$. Indeed, if all three local eigenvalues are in $\Lambda_{\beta}$ then by Theorem there exists a positive nonlocal eigenvalue $\nu\in(\lambda_{0},\lambda_{1})$. Otherwise at least one of those local eigenvalues is also a nonlocal eigenvalue (belongs to $\Lambda_{0}$) and hence the proof is finished. ∎
@@ 6.2. Direction of branches (sub- and supercritical pitchfork)
We differentiate $\mathcal{G}(s)$ with respect to $s$ up to third order. The derivatives of $G$ with respect to $\chi$ and $\alpha$ are given in Appendix B.
@@ 6.3. Global continuation and turning points
Since $\kappa_{1}$ is the smallest bifurcation point, the branch cannot intersect the trivial branch for $\kappa\in(0,\kappa_{1})$; this corresponds to the vertical line $A_{1}$ in Figure 6.1. Furthermore, Theorem 6.5 guarantees the uniqueness of the constant solution for $\kappa<\kappa_{\min}$, which forms the boundary $A_{2}$.

### v2

@@ 4 School of Mathematics and Physics, University of Queensland, Australia
This work was supported by the German Research Foundation (DFG) under Germany’s Excellence Strategy (EXC 2181/1 – 390900948, Heidelberg STRUCTURES Excellence Cluster) and through the Collaborative Research Center 1324 (SFB 1324, Project B6). Additional funding was provided by the European Research Council (ERC) under the Synergy Grant (PEPS, No. 10107178), and by the German Academic Exchange Service (DAAD) within the Project-Based Personnel Exchange Programme (PPP, Project No. 57820507).
@@ 1. Introduction
By deriving a reduced one-dimensional model that preserves the essential mechanochemical feedback while simplifying the geometric setting (Section 3, Appendix A), we obtain, for the first time, a framework amenable to rigorous analytical investigation of the dynamics. We establish the existence of spatially heterogeneous steady states for sufficiently small diffusion coefficients and prove uniqueness of the homogeneous equilibrium for large diffusion, deriving explicit bounds on the corresponding critical values by variational methods (Section 4). This analysis explains the mechanism of symmetry breaking, demonstrating how patterned states bifurcate from the homogeneous configuration. We then investigate the pattern selection problem (Section 5) and show that only unimodal stationary patterns are linearly stable, while all multimodal solutions are unstable, revealing a fundamental distinction from classical Turing systems and curvature-based mechanochemical models [38, 31]. Finally, we identify both subcritical and supercritical pitchfork bifurcations and reveal additional fold bifurcations responsible for the emergence of bistable regimes (Section 6). Throughout, the theoretical results are complemented by numerical simulations, elucidating the interplay between nonlocality, mechanical feedback, and nonlinear kinetics, and providing a mathematically grounded framework for experimental exploration.
@@ 2. Main results
We start by formulating main theoretical results of this paper. We focus the mechanochemical pattern formation model describing the concentration of the diffusive signaling molecules $u:=u(x,t)$, given by
@@
with the coefficients $D>0$ and $\kappa>0$, supplemented with the periodic boundary conditions $u(0,t)=u(1,t)$ and $\partial_{x}u(0,t)=\partial_{x}u(1,t)$ for all $t\geqslant 0$. This model arises from a one-dimensional formulation of a hyperelastic material governed by the Saint-Venant–Kirchhoff constitutive law. A detailed derivation is provided in Section 3 and Appendix A. Numerical simulations of the model (Figure 2.1) demonstrate its capacity to produce stable unimodal patterns.
@@
    $\Gamma_{n}(s)=\big(\bar{\varphi}_{n}(s),\kappa_{n}(s)\big),$
    admitting the following asymptotic expansion as $s\to 0$,
    (2.8) $\bar{\varphi}_{n}(s)=\kappa_{n}+s\sqrt{2}\cos(2\pi nx)+\mathcal{O}(s^{2}),\ \kappa_{n}(s)=\kappa_{n}+C_{n}\,s^{2}+\mathcal{O}(s^{3}),$
    where $C_{n}=\frac{1}{4}-\frac{1}{16Dn^{2}\pi^{2}}+2Dn^{2}\pi^{2}$.
@@ 3. Model formulation
Figure 3.1. Schematic representation of the elastic deformation framework for the Hydra spheroid model. Left: Circular reference configuration $\mathcal{S}_{0}$ parametrized by arc-length $x\in[0,L]$. Right: Deformed circular configuration $\mathcal{S}$ in which material points have arc-length position $s(x)\in[0,\ell]$.
We reformulate the mechanochemical patterning model for Hydra spheroids introduced in [41] in two spatial dimensions in which the spheroid is described as a closed curve. Starting from the model derivation and formulation in 3D using index notation, one can specialise the notation in 2D as it was done in the supplementary material of [41]. Alternatively, we derive the model directly, using the same modelling arguments but without using index notation, in the Appendix A.
@@
Ultimately, we derive a reduced spatially one-dimensional formulation of the original model in [41] that retains the essential coupling between mechanics and chemistry while simplifying the geometric structure. In this setting, the Hydra spheroid is represented as an elastic closed curve, providing a minimal framework to analyze the core mechanochemical feedback mechanisms, see Figure 3.1.
The stress-free reference configuration, denoted by $\mathcal{S}_{0}$, is assumed to be a circle of fixed circumference $L>0$. It is parametrised by $x\in[0,L]$ which denotes the arc-length associated to a given point in $\mathcal{S}_{0}$. We prescribe a fixed enclosed area $A_{0}>\operatorname{Area}(\mathcal{S}_{0})$ and consider deformations of the reference shape which enclose the area $A_{0}$. In Appendix A we show that, as a consequence of stress balance and the Law of Laplace, the deformed spheroid is always a circle of the given circumference $\ell>L$ and that no other shapes are possible.
The scalar function $s(x):[0,L]\mapsto[0,l]$, $s^{\prime}>0$, represents the arc-length position in the deformed spheroid $\mathcal{S}$ of the material point with arc-length position $x$ in $\mathcal{S}_{0}$. As the deformed spheroid is circular, any deformation is by shifting material points tangentially within the hydra tissue. The full mechanochemical system written in terms of the morphogen concentration $u=u(t,x)$ and the strain $\varepsilon(t,x)$ reads
@@
(3.1) $\left\{\begin{aligned} &\begin{aligned} \partial_{t}u&=\kappa\,\varepsilon-\alpha u+D\,\partial_{xx}u,\\ 0&=\partial_{x}\big(E(u)\,\varepsilon\big),\end{aligned}&&x\in[0,L],\quad t\geq 0,\\ &l-L=\int_{0}^{L}\varepsilon\,dx,&&t\geq 0.\end{aligned}\right$.
This system of equations combines stress balance where $E(u)>0$ is the elastic modulus given by a monotonically decreasing function of $u$. This models increased tissue elasticity (reduced stiffness) in response to higher expression of morphogen. This implies that solutions are characterised by larger strain where the morphogen concentration is high.
This ”mechanical” equation is coupled to a ”chemical” reaction diffusion equation for the morphogen concentration. Here, $D$ denotes diffusivity, $\alpha$ is the spontaneous disassociation rate and $\kappa>0$ is a given parameter associated to the expression of morphogen in response to tissue strain $\varepsilon$. Up-regulation of morphogen concentration in response to tissue strain complements increased strain where morphogen is over-expressed according to the mechanical part of the model. Together, these two components form a positive local feedback loop.
@@
Note that following infinitesimal stress theory, we write the strain as $\varepsilon=s^{\prime}-1$ which motivates the third equation in (3.1) and expresses conservation of total strain, effectively introducing mechanical long-range inhibition, which acts alongside the positive feedback between local strain and morphogen concentration. This combination implements the classical local activation - long-range inhibition (LALI) mechanism.
@@
From the momentum balance equation in (3.1), the stress depends only on time $t$, that is,
@@
(3.6) $\varepsilon(t,x)=\frac{\sigma(t)}{E(u(t,x))}=\frac{l-L}{E(u(t,x))\int_{0}^{L}\frac{1}{E(u(t,y))}\,dy}.$
Using the closed-form expression for the strain, the morphogen equation in (3.1) can be written as
(3.7) $\partial_{t}u(t,x)=\frac{\kappa(l-L)}{E(u(t,x))\int_{0}^{L}\frac{1}{E(u(t,y))}\,dy}-\alpha u+D\,\partial_{xx}u\,.$
@@ 4.1. Variational formulation
There exist $\{f_{k}\}_{k=0}^{\infty}$ an orthonormal basis of $L^{2}(\mathbb{T})$ and an orthogonal basis of $W^{1,2}(\mathbb{T})$ and a set of eigenvalues $\{\mu_{k}\}_{k=0}^{\infty}\subseteq[0,\;\infty)$, such that $(f_{k},\mu_{k})$ is the eigenpair of the operator $\Delta$
@@ 5.3. Relation between eigenvalues
We must account for the terms where $\beta_{n}=0$ as they vanish in the series (5.8). These sets are not necessarily disjoint, i.e. there can exists a double eigenvalue $\lambda_{n}=\lambda_{n+1}$ and $\lambda_{n}\in\Lambda_{0}$ while $\lambda_{n+1}\in\Lambda_{\beta}$. Without loss of generality we assume that the elements of $\Lambda_{0},\Lambda_{\beta}$ are in monotonically decreasing order
@@ 5.4. Stability of steady states
which is exactly the local problem (5.6) with $\lambda=0$. Since $U\in W^{2,2}(\mathbb{T})$, all $U$, $U_{x}$ and $U_{xx}$ satisfy periodic boundary conditions. Hence, the local problem has a zero eigenvalue with corresponding eigenfunction with at least $3$ zeros on the interval $[0,1)$. Proposition 5.3 implies that $\lambda_{3}\geqslant 0$ as the eigenfunctions $\psi_{0}$, $\psi_{1}$, $\psi_{2}$ have at most two zeros on the interval $[0,1)$. Therefore there are at least three strictly positive local eigenvalues $\lambda_{0}>\lambda_{1}\geqslant\lambda_{2}>\lambda_{3}\geq 0$. This is already sufficient to obtain a positive nonlocal eigenvalue $\nu$. Indeed, if all three local eigenvalues are in $\Lambda_{\beta}$ then by Theorem 5.5 there exists a positive nonlocal eigenvalue $\nu\in(\lambda_{0},\lambda_{1})$. Otherwise at least one of those local eigenvalues is also a nonlocal eigenvalue (belongs to $\Lambda_{0}$) and hence the proof is finished. ∎
@@ 6.2. Direction of branches (sub- and supercritical pitchfork)
We differentiate $\mathcal{G}(s)$ with respect to $s$ up to third order. The derivatives of $G$ with respect to $\chi$ and $\alpha$ are given in Appendix C.
@@ 6.3. Global continuation and turning points
Since $\kappa_{1}$ is the smallest bifurcation point, the branch cannot intersect the trivial branch for $\kappa\in(0,\kappa_{1})$; this corresponds to the horizontal line $A_{1}$ in Figure 6.1. Furthermore, Theorem 6.5 guarantees the uniqueness of the constant solution for $\kappa<\kappa_{\min}$, which forms the boundary $A_{2}$.